OpenStudy (anonymous):

Can someone tell me if I did this correctly? Simplify the following product of complex numbers. (-10 + 3i) * (-2) = (-20 + -6i )= -26i ???

OpenStudy (anonymous):

(-10 + 3i) * (-2)= (20-6i )= 14i

OpenStudy (anonymous):

yes the second one is correct.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

(20-6i ) this is as simplify as you can get it to be. Reason being both are not like terms so you can't subtract or add each other.

OpenStudy (anonymous):

ok thanks ryan, that does make sense :)